# MultiSensor-Home: A Wide-area Multi-modal Multi-view Dataset for Action Recognition and Transformer-based Sensor Fusion

### v1

@@ IV-A2 Overview of MultiTSF
To achieve robust performance in multi-modal multi-view action recognition, the proposed method integrates several key components, as illustrated in Figure 4. The Multi-modal Feature Extraction module utilizes the Audio Spectrogram Transformer (AST) [7] and Vision Transformer (ViT) [5] to extract discriminative features from audio and visual streams. The Human Detection Module employs the You Only Look Once (YOLOv10) model [31] to detect human presence and generate pseudo-ground-truth labels, guiding the model to effectively learn human activity within the visual features. To capture temporal dependencies and integrate multi-view spatiotemporal features, the Temporal Modeling and Transformer-based Fusion component leverages a Transformer-based attention mechanism. Finally, the Learning Objectives module optimizes the framework using frame-level, sequence-level, and human loss functions, enhancing spatial and temporal understanding for accurate action recognition.
@@ IV-B Multi-modal Feature Extraction
In this module, we employ a shared audio encoder and visual encoder to process audio and visual inputs. These shared encoders use the same model parameters for all input views, ensuring consistent and efficient feature extraction.
@@ IV-B1 Shared Audio Encoder
Each raw audio signal $A_{i}$ from the $i$-th view ($i\in\{1,2,\dots,N\}$) is transformed into a log-mel spectrogram, which captures the essential temporal and frequency characteristics of the audio. These spectrograms are then processed using a shared AST model [7] to extract discriminative features. The AST divides the spectrogram into overlapping patches, projects these patches into embeddings, and processes the sequence of embeddings through Transformer layers. The resulting audio features are represented as:
$F_{i}^{\mathsf{A}}=f_{a}(A_{i}),\quad F_{i}^{\mathsf{A}}\in\mathbb{R}^{T\times D_{\mathsf{A}}},$ (1)
where $f_{a}(\cdot)$ represents the AST model operation, $D_{\mathsf{A}}$ denotes the dimensionality of the extracted audio features.
@@ IV-B2 Shared Visual Encoder
To extract spatial features from each video frame in each video $V_{i}$ ($i\in\{1,2,\dots,N\}$), we use a shared ViT model [5]. The ViT splits each video frame into non-overlapping patches of size $P\times P$, flattens these patches, and projects them into embeddings. These embeddings are then processed by Transformer layers to capture spatial dependencies. The visual features for the $i$-th view are represented as:
$F_{i}^{\mathsf{V}}=f_{v}(V_{i}),\quad F_{i}^{\mathsf{V}}\in\mathbb{R}^{T\times D_{\mathsf{V}}},$ (2)
where $f_{v}(\cdot)$ represents the ViT model operation, where $D_{\mathsf{V}}$ is the dimensionality of the extracted visual features.
@@ IV-B3 Audio-Visual Features
The extracted audio features $F_{i}^{\mathsf{A}}$ and visual features $F_{i}^{\mathsf{V}}$ are concatenated at the frame level to form combined audio-visual features. For each frame $t\in\{1,2,\dots,T\}$, the audio-visual feature is computed as:
$F_{i}^{\mathsf{AV}}(t)=[F_{i}^{\mathsf{A}}(t);F_{i}^{\mathsf{V}}(t)],$ (3)
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes concatenation along the feature dimension. The combined sequence of audio-visual features for the $i$-th view is represented as $F_{i}^{\mathsf{AV}}\in\mathbb{R}^{T\times D_{\mathsf{AV}}},$ where $D_{\mathsf{AV}}=D_{\mathsf{A}}+D_{\mathsf{V}}$. These fused features are subsequently passed to the temporal modeling and fusion stages for further processing.
@@ IV-C Human Detection Module
To enhance spatial feature learning, we introduce an Human Detection Module that detects human presence in video frames and generates pseudo-ground-truth labels for the Human Loss function ($\mathcal{L}_{\text{H}}$). This module operates independently of the main framework and processes visual data from multiple views. For each video $V_{i}$ from the $i$-th view ($i\in\{1,2,\dots,N\}$), the frames are passed through the object detection model (i.e., YOLOv10 model [31]). The YOLOv10 model outputs a binary indicator for each frame $t\in\{1,2,\dots,T\}$, denoted as:
$h_{i}(t)=\begin{cases}1,&\text{if a human is detected in frame }t,\\ 0,&\text{otherwise}.\end{cases}$ (4)
The binary outputs $h_{i}(t)$ for all frames serve as pseudo-ground truth labels for supervising the learning of human-related spatial features. These labels guide the model to focus on frames containing human activity and ignore irrelevant frames. The pseudo-ground-truth labels $h_{i}(t)$ are integrated into the model as supervision for learning relevant spatial features. During training, these labels are used in the Human Loss function ($\mathcal{L}_{\text{H}}$), which is detailed in the Learning Objectives section.
@@ IV-D1 Shared Temporal Encoder
We employ a Shared Temporal Encoder to capture temporal dependencies in the extracted audio-visual features. The input to the encoder is the concatenated audio-visual features $F_{i}^{\mathsf{AV}}$ from each view $i\in\{1,2,\dots,N\}$. The Shared Temporal Encoder uses a self-attention mechanism [28] to model the temporal relationships across frames within a view. Specifically, the self-attention mechanism computes the weighted interactions between frames:
@@
where $Q,K,V\in\mathbb{R}^{T\times d_{k}}$ are the query, key, and value matrices derived from $F_{i}^{\mathsf{AV}}$, and $d_{k}$ is the dimensionality of the query and key spaces. The output of this encoder for each view $i$ is a set of temporally enhanced features:
$F_{i}^{\mathsf{T}}=\left[F_{i}^{\mathsf{T}}(1),F_{i}^{\mathsf{T}}(2),\dots,F_{i}^{\mathsf{T}}(T)\right]\in\mathbb{R}^{T\times D_{\mathsf{T}}},$ (6)
@@ IV-D2 Transfomer-based Sensor Fusion
After extracting temporal features $F_{i}^{\mathsf{T}}$ from all $N$ views, the Transformer-based Sensor Fusion mechanism combines information from multiple views to generate a unified representation. For each frame $t\in\{1,2,\dots,T\}$, the temporal features from all $N$ views are processed to capture inter-view relationships. Specifically, for a given frame $t$, the temporal features from the $N$ views are represented as:
@@
These features are input to the self-attention to model the importance and relationships between views for the same frame $t$. This mechanism allows the model to assign dynamic importance to different views based on their contributions to the frame’s action recognition. The output is a fused feature for the frame $F^{\mathsf{Fusion}}(t)\in\mathbb{R}^{D_{\mathsf{Fusion}}},$ where $D_{\mathsf{Fusion}}$ is the dimensionality of the fused features. Repeating this process for all $T$ frames yields the final sequence of fused features:
@@ IV-E Learning Objectives
We design the learning objectives to optimize spatial, temporal, and inter-view features for effective action recognition. The Human Loss ($\mathcal{L}_{\text{H}}$) ensures focus on frames with human activity. The Frame Loss ($\mathcal{L}_{\text{F}}$) and Sequence Loss ($\mathcal{L}_{\text{S}}$) address class imbalance at the frame-level and sequence-level, respectively. The total loss function is computed as:
@@
IV-E1 Human Loss ($\mathcal{L}_{\text{H}}$)
This loss ensures that the model effectively learns to identify whether a frame contains human activity across $N$ views. For each view $i\in\{1,\dots,N\}$, the Shared Visual Encoder extracts frame-level features. These features are processed through a linear layer to predict the probability $\hat{h}_{i}(t)$ of human presence in frame $t$. The Binary Cross Entropy loss is used to measure the discrepancy between the predicted probabilities $\hat{h}_{i}(t)$ and the pseudo-ground-truth labels $h_{i}(t)$:
$\mathcal{L}_{\text{H}}=-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\Big{[}$ $h_{i}(t)\log\hat{h}_{i}(t)$ (10)
$+\big{(}1-h_{i}(t)\big{)}\log\big{(}1-\hat{h}_{i}(t)\big{)}\Big{]}.$
IV-E2 Frame Loss ($\mathcal{L}_{\text{F}}$)
This loss optimizes the model’s ability to predict frame-level action classes. To address class imbalance, we employ a two-way loss function [12], which combines sample-wise and class-wise components:
@@ IV-E2 Frame Loss ($\mathcal{L}_{\text{F}}$)
Sample-wise loss for frames discriminates between positive and negative classes:
$\small\mathcal{L}_{\text{F}}^{\mathsf{S}}=\frac{1}{T}\sum_{t=1}^{T}\text{softplus}\left(\log\sum_{n|y_{t}=0}e^{x_{\mathsf{S}_{n}}}+\gamma_{s}\log\sum_{p|y_{t}=1}e^{-\frac{x_{\mathsf{S}_{p}}}{\gamma}}\right),$ (12)
where $y_{t}$ represents the ground-truth label for the $t$-th frame, $x_{\mathsf{S}_{n}}$ and $x_{\mathsf{S}_{p}}$ are the logits for negative and positive classes, respectively, and $\gamma_{s}$ is a temperature parameter. The $\text{softplus}(\cdot)=\log(1+\text{exp}(\cdot))$ is a smooth approximation to the Rectified Linear Unit (ReLU) function.
Class-wise loss for frames addresses intra-class variations:
$\small\mathcal{L}_{\text{F}}^{\mathsf{C}}=\frac{1}{C}\sum_{c=1}^{C}\text{softplus}\left(\log\sum_{n|y_{c}=0}e^{x_{\mathsf{C}_{n}}}+\gamma_{c}\log\sum_{p|y_{c}=1}e^{-\frac{x_{\mathsf{C}_{p}}}{\gamma}}\right),$ (13)
where $y_{c}$ represents the ground-truth label for class $c$, $x_{\mathsf{C}_{n}}$ and $x_{\mathsf{C}_{p}}$ are the logits for negative and positive samples within class $c$, and $\gamma_{c}$ is a temperature parameter.
IV-E3 Sequence Loss ($\mathcal{L}_{\text{S}}$)
This loss optimizes sequence-level predictions by aggregating temporal features across $T$ frames. The fused sequence features are passed through a classification head, and the loss is calculated using the two-way loss [12]:
@@ IV-E3 Sequence Loss ($\mathcal{L}_{\text{S}}$)
where the sample-wise and class-wise loss components are computed similarly to Frame loss but applied to sequence-level predictions, $\alpha_{2}$ is a balancing parameter.
@@ V Performance Evaluation
In this section, we extensively evaluate MultiTSF on the MultiSensor-Home dataset and the MM-Office dataset [35]. The analysis includes detailed quantitative, qualitative, and ablation studies to assess the robustness of the proposed method across diverse datasets and to understand the contribution of individual components within MultiTSF.
@@ V-A1 Data Preparation
We evaluate the proposed method using the MultiSensor-Home dataset and MM-Office dataset [35]. We apply the iterative stratification strategy [22] to divide the data into training and test subsets, ensuring a balanced representation of all classes in these subsets. The MultiSensor-Home dataset is split in a 70:30 ratio [^2] Details of the train/test splits are in the Supplemental Material., while the MM-Office dataset follows the splitting strategy outlined in [18].
For the experiments, we extract a fixed number of $T$ synchronized frames, where visual frames are sampled uniformly from the video at a fixed frame rate (e.g., 2.5 FPS) to ensure consistent temporal spacing, and corresponding audio segments are aligned based on the exact timestamps of these frames. During the training phase, we generate a sequence of frame indices using uniform sampling with random perturbations, ensuring that the sequence spans the entire video while maintaining a fixed length of $T$. This approach serves as a data augmentation technique, enhancing the robustness of the model by introducing variability during training. For testing, we apply uniform sampling without perturbation to ensure consistency across all test runs.
@@ V-A3 Comparison Methods
We compare the proposed MultiTSF method with several state-of-the-art approaches for multi-modal multi-view and video-based action recognition. Specifically, for multi-modal multi-view methods, we include MultiTrans [35] and MultiASL [18]. For video-based action recognition, we evaluate against TimeSformer [3] and ViViT [1], which are Transformer-based architectures that capture spatiotemporal relationships, and X-CLIP [16], which extends CLIP [21] to video data.
V-A4 Models & Hyperparameters
We implement MultiTSF as detailed in Section IV [^3] The source code is provided in Supplemental Material.. For MultiSensor-Home and MM-Office datasets, we use $T=70$ and $T=50$, respectively, sampled at 2.5 FPS, based on the average video lengths in each dataset. The Shared Audio Encoder extracts 128-dimensional log-mel filterbank features using the AST [7], while the Shared Video Encoder processes frames with a resolution of $224\times 224$ pixels using ViT [5]. Shared Temporal Encoder uses two-layer Transformer encoder with four attention heads of 128 dimensions. Transformer-based Sensor Fusion uses a single layer with four attention heads of 128 dimensions. The hyperparameters controlling the relative importance of each loss term are set to 1 for simplicity. Optimization is performed using Adam [11] with an initial learning rate of $10^{-4}$, a weight decay of $5.0\times 10^{-4}$, and a batch size of 12 for 300 epochs. The learning rate is scheduled using Cosine Annealing [15] to adaptively decay over training. All experiments are conducted on a machine with four Tesla V100-PCIE-32GB GPUs.
TABLE III: Comparison of MultiTSF with other methods on the MultiSensor-Home and MM-Office datasets. The best and second-best results are highlighted in bold and underlined text, respectively.
@@ V-A4 Models & Hyperparameters
Multi-modal (Visual + Audio)
MultiTrans [35] 59.65 77.60 61.40 78.07
MultiASL [18] 58.58 77.43 73.81 85.38
MultiTSF (Proposed) 64.48 87.91 76.12 91.45

### v3

@@ IV-A2 Overview of MultiTSF
To achieve robust performance in multi-modal multi-view action recognition, the proposed method integrates several key components, as illustrated in Figure 4. The Multi-modal Feature Extraction module extracts discriminative features from audio and visual streams. The Human Detection Module employs an object detection model to detect frames with human presence and generate pseudo-ground-truth labels, guiding the model to effectively learn human activity within the visual features. To capture temporal dependencies and integrate multi-view spatiotemporal features, the Temporal Modeling and Transformer-based Fusion component leverages a Transformer-based attention mechanism. Finally, the Learning Objectives module optimizes the framework using frame-level, sequence-level, and human loss functions, enhancing spatial and temporal understanding for accurate action recognition.
@@ IV-B Multi-modal Feature Extraction
In this module, we employ a Shared Audio Encoder and a Shared Visual Encoder to process audio and visual input data. These encoders use the same model parameters for all input views, ensuring consistent and efficient feature extraction.
@@ IV-B1 Shared Audio Encoder
Each raw audio signal $A_{n}$ from the $n$-th view ($n\in\{1,2,\dots,N\}$) is transformed into a log-mel spectrogram, which captures the temporal and frequency characteristics of the audio. These spectrograms are then processed using a shared Audio Spectrogram Transformer (AST) model [7] to extract discriminative features. AST divides the spectrogram into overlapping patches, projects them into embeddings, and processes the sequence of embeddings through Transformer layers. The resulting audio features are represented as:
$F_{n}^{\mathsf{A}}=f_{a}(A_{n}),\quad F_{n}^{\mathsf{A}}\in\mathbb{R}^{T\times D_{\mathsf{A}}},$ (1)
where $f_{a}(\cdot)$ represents the AST model operation, and $D_{\mathsf{A}}$ is the dimensionality of the extracted audio features.
@@ IV-B2 Shared Visual Encoder
To extract spatial features from each video frame in each video $V_{n}$, we use a shared Vision Transformer (ViT) model [5]. ViT splits each video frame into non-overlapping patches of size $P\times P$ [pixels], flattens these patches, and projects them into embeddings. These embeddings are then processed by Transformer layers to capture spatial dependencies. The visual features for the $n$-th view are represented as:
$F_{n}^{\mathsf{V}}=f_{v}(V_{n}),\quad F_{n}^{\mathsf{V}}\in\mathbb{R}^{T\times D_{\mathsf{V}}},$ (2)
where $f_{v}(\cdot)$ represents the ViT model operation, and $D_{\mathsf{V}}$ is the dimensionality of the extracted visual features.
@@ IV-B3 Audio-Visual Features
The extracted audio features $F_{n}^{\mathsf{A}}$ and visual features $F_{n}^{\mathsf{V}}$ are concatenated at the frame level to form combined audio-visual features. For each frame $t\in\{1,2,\dots,T\}$, the audio-visual feature is computed as:
$F_{n}^{\mathsf{AV}}(t)=[F_{n}^{\mathsf{A}}(t);F_{n}^{\mathsf{V}}(t)],$ (3)
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes concatenation along the feature dimension. The combined sequence of audio-visual features for the $n$-th view is represented as $F_{n}^{\mathsf{AV}}\in\mathbb{R}^{T\times D_{\mathsf{AV}}},$ where $D_{\mathsf{AV}}=D_{\mathsf{A}}+D_{\mathsf{V}}$. These fused features are subsequently passed to the temporal modeling and fusion stages for further processing.
@@ IV-C Human Detection Module
To enhance spatial feature learning, we introduce a Human Detection Module that detects human presence in video frames and generates pseudo-ground-truth labels for the Human Loss function. This module operates independently of the main framework and processes visual data from multiple views. For each video $V_{n}$ from the $n$-th view ($n\in\{1,2,\dots,N\}$), the frames are passed through the Object Detection Model (i.e., You Only Look Once (YOLO) v10 model [31]), which outputs a binary indicator for each frame $t\in\{1,2,\dots,T\}$ as:
$h_{n}(t)=\begin{cases}1,&\text{if a human is detected in frame }t,\\ 0,&\text{otherwise}.\end{cases}$ (4)
The binary outputs $h_{n}(t)$ for all frames serve as pseudo-ground-truth labels to supervise the learning of human-related spatial features. These labels guide the model to focus on frames with human activity and ignore irrelevant frames. During training, they are used in the Human Loss function, which is detailed in the Learning Objectives section.
@@ IV-D1 Shared Temporal Encoder
We employ a Shared Temporal Encoder to capture temporal dependencies in the extracted audio-visual features. The input to the encoder is the concatenated audio-visual features $F_{n}^{\mathsf{AV}}$ from each view $n\in\{1,2,\dots,N\}$. The Shared Temporal Encoder uses a self-attention mechanism [28] to model the temporal relationships across frames within a view. Specifically, the self-attention mechanism computes the weighted interactions between frames as:
@@
where $Q,K,V\in\mathbb{R}^{T\times d_{k}}$ are the query, key, and value matrices derived from $F_{n}^{\mathsf{AV}}$, and $d_{k}$ is the dimensionality of the query and key spaces. The output of this encoder for each view $n$ is a set of temporally enhanced features as:
$F_{n}^{\mathsf{T}}=\left[F_{n}^{\mathsf{T}}(1),F_{n}^{\mathsf{T}}(2),\dots,F_{n}^{\mathsf{T}}(T)\right]\in\mathbb{R}^{T\times D_{\mathsf{T}}},$ (6)
@@ IV-D2 Transfomer-based Sensor Fusion
After extracting temporal features $F_{n}^{\mathsf{T}}$ from all $N$ views, the Transformer-based Sensor Fusion mechanism combines information from multiple views to generate a unified representation. For each frame $t\in\{1,2,\dots,T\}$, the temporal features from all $N$ views are processed to capture inter-view relationships. Specifically, for a given frame $t$, the temporal features from the $N$ views are represented as:
@@
These features are input to the self-attention mechanism to model the importance and relationships between views for the same frame $t$. This mechanism allows the model to assign dynamic importance to different views based on their contributions to the frame’s action recognition. The output is a fused feature for the frame $F^{\mathsf{Fusion}}(t)\in\mathbb{R}^{D_{\mathsf{Fusion}}},$ where $D_{\mathsf{Fusion}}$ is the dimensionality of the fused features. Repeating this process for all $T$ frames yields the final sequence of fused features as:
@@ IV-E Learning Objectives
We design the learning objectives to optimize spatial, temporal, and inter-view features for effective action recognition. Human Loss $\mathcal{L}_{\text{H}}$ ensures focus on frames with human activity. Frame Loss $\mathcal{L}_{\text{F}}$ and Sequence Loss $\mathcal{L}_{\text{S}}$ address class imbalance at the frame-level and sequence-level, respectively. The total loss function is computed as:
@@
IV-E1 Human Loss $\mathcal{L}_{\mathrm{H}}$
This loss ensures that the model effectively learns to identify whether a frame contains human activity across $N$ views. For each view $n\in\{1,\dots,N\}$, the Shared Visual Encoder extracts frame-level features. These features are processed through a linear layer to predict the probability $\hat{h}_{n}(t)$ of human presence in frame $t\in\{1,2,\dots,T\}$. The Binary Cross Entropy loss is used to measure the discrepancy between the predicted probabilities $\hat{h}_{n}(t)$ and the pseudo-ground-truth labels $h_{n}(t)$ as:
$\mathcal{L}_{\text{H}}=-\frac{1}{NT}\sum_{n=1}^{N}\sum_{t=1}^{T}\Big{[}$ $h_{n}(t)\log\hat{h}_{n}(t)$ (10)
$+\big{(}1-h_{n}(t)\big{)}\log\big{(}1-\hat{h}_{n}(t)\big{)}\Big{]}.$
IV-E2 Frame Loss $\mathcal{L}_{\mathrm{F}}$
This loss optimizes the model’s ability to predict frame-level action classes. To address class imbalance, we employ a two-way loss function [12], which combines sample-wise and class-wise components as:
@@ IV-E2 Frame Loss $\mathcal{L}_{\mathrm{F}}$
Here, the sample-wise loss $\mathcal{L}_{\text{F}}^{\mathsf{S}}$ for frames discriminates between positive and negative classes as:
$\small\mathcal{L}_{\text{F}}^{\mathsf{S}}=\frac{1}{T}\sum_{t=1}^{T}\text{softplus}\left(\log\sum_{n\in\mathcal{N}_{t}}e^{x_{\mathsf{S}_{n}}}+\gamma_{s}\log\sum_{p\in\mathcal{P}_{t}}e^{-\frac{x_{\mathsf{S}_{p}}}{\gamma_{s}}}\right),$ (12)
where $\mathcal{N}_{t}$ and $\mathcal{P}_{t}$ denote the index sets of negative and positive samples for frame $t$, respectively; $x_{\mathsf{S}_{n}}$ and $x_{\mathsf{S}_{p}}$ are the logits corresponding to negative and positive samples; and $\gamma_{s}$ is a temperature parameter. The function $\text{softplus}(\cdot)=\log(1+\exp(\cdot))$ serves as a smooth approximation to the Rectified Linear Unit (ReLU) function.
Meanwhile, the class-wise loss $\mathcal{L}_{\text{F}}^{\mathsf{C}}$ addresses intra-class variations across frames as:
$\small\mathcal{L}_{\text{F}}^{\mathsf{C}}=\frac{1}{C}\sum_{c=1}^{C}\text{softplus}\left(\log\sum_{n\in\mathcal{N}_{c}}e^{x_{\mathsf{C}_{n}}}+\gamma_{c}\log\sum_{p\in\mathcal{P}_{c}}e^{-\frac{x_{\mathsf{C}_{p}}}{\gamma_{c}}}\right),$ (13)
where $\mathcal{N}_{c}$ and $\mathcal{P}_{c}$ represent the index sets of negative and positive samples within class $c$, respectively; $x_{\mathsf{C}_{n}}$ and $x_{\mathsf{C}_{p}}$ are their corresponding logits, and $\gamma_{c}$ is a temperature parameter.
IV-E3 Sequence Loss $\mathcal{L}_{\mathrm{S}}$
This loss optimizes sequence-level predictions by aggregating temporal features across $T$ frames. The fused sequence features are passed through a classification head, and the loss is calculated using the two-way loss [12] as:
@@ IV-E3 Sequence Loss $\mathcal{L}_{\mathrm{S}}$
where Sample-wise loss $\mathcal{L}_{\text{S}}^{\mathsf{S}}$ and Class-wise loss $\mathcal{L}_{\text{S}}^{\mathsf{C}}$ are computed similarly to Frame Loss $\mathcal{L}_{\text{F}}$ but applied to sequence-level predictions, and $\alpha_{2}$ is a balancing parameter.
@@ V-A1 Data Preparation
We evaluate the proposed method using the proposed MultiSensor-Home and the existing MM-Office [35] datasets. We apply the iterative stratification strategy [22] to divide the data into training and test subsets, ensuring a balanced representation of all classes in these subsets. The MultiSensor-Home dataset is split in a 70:30 ratio, while the MM-Office dataset follows the splitting strategy outlined in [18].
For the experiments, we extract a fixed number of $T$ synchronized frames, where visual frames are sampled uniformly from the video at a fixed frame rate (i.e., 2.5 FPS) to ensure consistent temporal spacing, and corresponding audio segments are aligned based on the exact timestamps of these frames. During the training phase, we generate a sequence of frame indices using uniform sampling with random perturbations, ensuring that the sequence spans the entire video while maintaining a fixed length of $T$. This approach serves as a data augmentation technique, enhancing the robustness of the model by introducing variability during training. For testing, we apply uniform sampling without perturbation to ensure consistency across all test runs.
@@ V-A3 Comparison Methods
We compare the proposed MultiTSF method with several state-of-the-art approaches for multi-modal multi-view and video-based action recognition. Specifically, for multi-modal multi-view methods, we include MultiTrans [35] and MultiASL [18]. For video-based action recognition, we evaluate against TimeSformer [3] and Video Vision Transformer (ViViT) [1], which are Transformer-based architectures that capture spatiotemporal relationships, and X-CLIP [16], which extends Contrastive Language-Image Pretraining (CLIP) [21] to video data.
TABLE III: Comparison of the proposed MultiTSF method with other methods on the MultiSensor-Home and MM-Office datasets. The best and second-best results are highlighted in bold and underlined text, respectively.
@@
Multi-modal (Audio + Visual)
MultiTrans [35] 59.65 77.60 61.40 78.07
MultiASL [18] 58.58 77.43 73.81 85.38
MultiTSF (Proposed) 64.48 87.91 76.12 91.45
@@
V-A4 Models & Hyperparameters
We implement the proposed MultiTSF as detailed in Section IV [^1] The source code is available at https://github.com/thanhhff/MultiTSF.. For MultiSensor-Home and MM-Office datasets, we use $T=70$ and $T=50$, respectively, sampled at 2.5 FPS, based on the average video lengths in each dataset. The Shared Audio Encoder extracts 128-dimensional log-mel filterbank features using AST [7], while the Shared Video Encoder processes frames with a resolution of $224\times 224$ pixels using ViT [5]. The Shared Temporal Encoder uses a two-layer Transformer encoder with four attention heads of 128 dimensions. Transformer-based Sensor Fusion uses a single layer with four attention heads of 128 dimensions. The hyperparameters controlling the relative importance of each loss term are set to 1 for simplicity. Optimization is performed using Adaptive moment estimation (Adam) [11] with an initial learning rate of $10^{-4}$, a weight decay of $5.0\times 10^{-4}$, and a batch size of 12 for 300 epochs. The learning rate is scheduled using Cosine Annealing [15] to adaptively decay over training. All experiments are conducted on a machine with four Tesla V100-PCIE-32GB GPUs.